# SSD4Rec: A Structured State Space Duality Model for Efficient Sequential Recommendation

### v1

@@ Abstract.
Sequential recommendation methods are crucial in modern recommender systems for their remarkable capability to understand a user’s changing interests based on past interactions. However, a significant challenge faced by current methods (e.g., RNN- or Transformer-based models) is to effectively and efficiently capture users’ preferences by modeling long behavior sequences, which impedes their various applications like short video platforms where user interactions are numerous. Recently, an emerging architecture named Mamba, built on state space models (SSM) with efficient hardware-aware designs, has showcased the tremendous potential for sequence modeling, presenting a compelling avenue for addressing the challenge effectively. Inspired by this, we propose a novel generic and efficient sequential recommendation backbone, SSD4Rec, which explores the seamless adaptation of Mamba for sequential recommendations. Specifically, SSD4Rec marks the variable- and long-length item sequences with sequence registers and processes the item representations with bidirectional Structured State Space Duality (SSD) blocks. This not only allows for hardware-aware matrix multiplication but also empowers outstanding capabilities in variable-length and long-range sequence modeling. Extensive evaluations on four benchmark datasets demonstrate that the proposed model achieves state-of-the-art performance while maintaining near-linear scalability with user sequence length. https://github.com/ZhangYifeng1995/SSD4Rec.
Sequential Recommendation, Recommender Systems, Mamba, State Space Model (SSM), State Space Duality (SSD), Transformer.
* Corresponding author: Wenqi Fan, Department of Computing, and Department of Management and Marketing, The Hong Kong Polytechnic University.
[…]
journalyear: 2024ccs: Information systems Recommender systems
@@ 1. Introduction
However, directly applying Mamba to sequential recommendation remains highly challenging, primarily due to two key hurdles. Due to varying length of users’ historical sequences towards items, most existing sequential modeling methods need padding or truncation operations for sequences with different lengths in recommender systems, resulting in excessive computational overhead (padding) or the potential loss of critical information that could enhance recommendations (truncation), as illustrated in Figure 2. Thus, the first challenge is how to model the dynamic users’ preference with variable- and lone-length historical interaction sequences. Meanwhile, naive Mamba-based models employ a left-to-right uni-directional modeling paradigm, which limits the power of hidden representations for items in historical sequences (Sun et al., 2019). Specifically, each item can only encode information from preceding items, making it difficult to capture precise user preferences and achieve optimal performance in sequential recommendations. This is due to the fact that users’ historical interactions might not follow a rigidly ordered sequence in practice. Thus, the second challenge is how to capture context from both preceding items and subsequent items within the users’ interaction history.
To address such challenges, in this paper, a Structured State-Space Duality-empowered recommendation framework (SSD4Rec) is proposed based on the advanced Mamba architecture to capture the dynamics underlying variable- and long-length user interaction sequences effectively and efficiently. A novel strategy is introduced to encode the user’s sequential interactions into a variable-length item sequence without requiring padding and truncation, thereby accommodating an unlimited input window. Meanwhile, a Bidirectional State Space Duality (Bi-SSD) block based on the Mamba architecture is designed to swallow the numerous item representations and predict corresponding user preferences for proper recommendations.
@@
  • We propose a novel strategy to encode the users’ historical behavior with varying lengths and a Bidirectional Structured State-Space Duality (Bi-SSD) block to capture contextual information for user behavior sequence modeling, effectively addressing the challenges of adapting Mamba for sequential recommendations.
  • We conducted extensive experiments on four commonly used datasets to empirically showcase the efficacy of SSD4Rec, including its superior recommendation performance and its scalability in predicting the preferences of users with variable- and long-length interaction histories. Specifically, SSD4Rec is 154.62% and 66.21% faster than the representative attention-based and SSM-based methods (i.e., SASRec (Kang and McAuley, 2018) and Mamba4Rec (Liu et al., 2024)), respectively, when training to extract features on user interactions at the length of 200.
@@ 2.1. Notations and Definitions
The primary goal of sequential recommendation is to predict the next interaction within a sequence. Let $\mathcal{U}=\{u_{1},u_{2},...,u_{|\mathcal{U}|}\}$ and $\mathcal{V}=\{v_{1},v_{2},...,v_{|\mathcal{V}}|\}$ be the sets of users and items, respectively. Each user $u_{k}\in\mathcal{U}$ is characterised by a chronologically ordered sequence $\mathcal{S}_{k}=[v^{u_{k}}_{1},...,v^{u_{k}}_{t},...,v^{u_{k}}_{L_{k}}]$, where $v^{u_{k}}_{t}\in\mathcal{V}$ represents the item that $u_{k}$ has interacted with at time step $t$, and $L_{k}>1$ denotes the interaction sequence length for user $u_{k}$, which may vary from user to user. Notably, for typical sequential recommendation methods, the input sequence length will be uniformed into a fixed hyper-parameter $L$ using the special padding token and truncation. Specifically, if $L_{k}<L$, the sequence is modified to $S_{k}^{L}=[pad,...,pad,v^{u_{k}}_{1},...,v^{u_{k}}_{L_{k}}]$, with the number of padding instances equal to $(L-L_{k})$. In contrast, if $L_{k}>L$, the sequence is truncated into $S_{k}^{L}=[v^{u_{k}}_{(L_{k}-L)},...,v^{u_{k}}_{L_{k}}]$. To represent items efficiently, it is common practice to use an embedding table $\mathcal{Q}=\{\mathbf{q}_{1},\mathbf{q}_{2},...,\mathbf{q}_{|\mathcal{V}|}\}$ for the item base $\mathcal{V}$, where $\mathbf{q}_{j}$ is a latent dense vector with dimensionality $d$ for item $v_{j}$.
@@
where $\hat{\mathbf{p}}_{k}$ denotes the preference representation of user $u_{k}$, and $r_{k}$ is a score distribution of user $u_{k}$ over all items in the item base $\mathcal{Q}$. This distribution is used to generate a ranked list of items representing the most likely interactions for user $u_{k}$ at step $L+1$.
@@ 2.2. An Overview of the Proposed SSD4Rec
Mamba models have recently demonstrated significant achievements across diverse domains, owing to their exceptional ability to capture intricate dynamics in long sequence modeling (Gu and Dao, 2023; Dao and Gu, 2024). By introducing advanced Structured State Space Models with hardware-aware expansions, Mamba facilitates attention-like modeling capabilities while maintaining a linear complexity in computing. Motivated by these advancements, this paper proposes to explore Mamba’s potential for effective and efficient sequential recommendations. Specifically, unlike the conventional sequential recommendation methods based on RNN or Transformer architecture, the proposed SSD4Rec is designed to handle variable-length item sequences seamlessly without requiring padding or truncation, thereby ensuring that the model receives complete information and retains essential knowledge. As shown in Figure 1, the proposed model consists of two key modules, namely variable-length item sequence construction and Bidirectional Structured State Space Duality (Bi-SSD) for user modeling. The first module aims to tackle the core challenge of managing variable-length sequences from diverse users within a certain batch. Meanwhile, the proposed block enables attention-like content-aware interaction modeling while upholding linear complexity in computation. Further elaboration on the proposed method will be provided in the upcoming sections.
@@
2.3. Variable-length Item Sequence Construction
@@
(3) $\mathcal{E}_{b}=[\underbrace{\mathcal{E}_{u_{1}},...,\mathcal{E}_{u_{k}},...,\mathcal{E}_{u_{K}}}_{\text{$b$-th batch with size $K$}}],$
where $\mathcal{E}_{u_{k}}\in\mathbb{R}^{L_{k}\times D}$ denotes the sequence of item embeddings $\{\mathbf{q}_{1}^{u_{k}},...,\mathbf{q}_{L_{k}}^{u_{k}}\}$ for user $u_{k}$ with the sequence length of $L_{k}$, and $\mathcal{E}_{b}\in\mathbb{R}^{1\times(\sum L)\times D}$ represents the integrated embedding sequence in the $b$-th batch. To delineate separate interactions, we incorporate segment registers based on user identifiers (IDs) with the integrated batch input. The registers can be written by
@@ 2.4. Bidirectional State Space Duality (Bi-SSD) for User Modeling
2.4.1. Structured State Space Duality
@@
(8) $\mathcal{F}=\textbf{SSM}(\mathbf{A},\mathbf{B},\mathbf{C})(\mathcal{X})$ $=(\mathbf{M}_{1}\mathcal{E}_{u_{1}},...,\mathbf{M}_{K}\mathcal{E}_{u_{K}}),$
@@
where $\mathcal{F}$ is the intermediate output of an SSD layer, which will be passed as the input sequence to the next SSD layer. $\mathbf{M}_{k}$ and $\mathcal{E}_{u_{k}}$ represent the structured matrix and item sequence for user $u_{k}$. By adopting this approach, SSD4Rec implements batch processing of user interaction sequences with varying lengths.
2.4.2. Block Design
By leveraging the state space duality (SSD) layer as the foundational component, we construct a network architecture tailored to address the sequential recommendation task, as illustrated in Figure 1. First, in addition to the standard sequential information concerning user behaviors, we propose integrating reverse insights from the interaction history by constructing a bidirectional architecture on top of the unidirectional SSD layer. In which, we introduce a weighted indicator $\beta$, to regulate the message integration between the forward and backward sequence modeling, expressed as
(10) $\mathcal{F}^{\prime}=\mathcal{F}_{\rm forward}+\beta*\mathcal{F}_{\rm backward},$
@@ 2.4.2. Block Design
where $\beta$ is a hyper-parameter to balance the contribution of backward sequence modeling. Empirically, $\beta$ should be set within the range of 0.1 to 0.3. This bidirectional design encourages the model to learn to reason in a reverse manner, potentially capturing more valuable patterns within the interaction histories. Notably, to conserve computational resources and GPU memory usage, we can utilize a shared SSD layer for the forward and backward modeling processes. Second, the utilization of SSD layers enables our model to facilitate hardware-aware computation, i.e., block-based matrix multiplication (Dao and Gu, 2024), thereby achieving attention-like content-aware modeling while preserving efficient learning. Moreover, by stacking multiple bidirectional SSD cells, the proposed model can achieve a deeper network structure with an increased number of learnable parameters. Layer-wise normalization and embedding dropout are further utilized between the stacked cells to improve the model’s robustness and stabilize the learning process. Finally, the proposed block predicts the representations of user preferences in a specific batch, which are used to calculate the ranking score distributions for Top-K sequential recommendations. Specifically, given the integrated input $\mathcal{X}=\{\mathcal{E}_{b},\mathcal{R}_{b}\}$ associated with $K$ users in the $b$-th batch, the proposed Bi-SSD layer outputs the representations of their user preferences $\mathcal{P}_{b}=\{\hat{\mathbf{p}}_{1},...,\hat{\mathbf{p}}_{k},...,\hat{\mathbf{p}}_{K}\}$, correspondingly. The process can be formulated as
(11) $\mathcal{P}_{b}=\textbf{Bi-SSD}(\mathcal{E}_{b},\mathcal{R}_{b}),$
where $\hat{\mathbf{p}}_{k}\in\mathbb{R}^{D}$ denotes the predicted representation for user $u_{k}$.

### v2

@@ Abstract.
Sequential recommendation methods are crucial in modern recommender systems for their remarkable capability to understand a user’s changing interests based on past interactions. However, a significant challenge faced by current methods (e.g., RNN- or Transformer-based models) is to effectively and efficiently capture users’ preferences by modeling long behavior sequences, which impedes their various applications like short video platforms where user interactions are numerous. Recently, an emerging architecture named Mamba, built on state space models (SSM) with efficient hardware-aware designs, has showcased the tremendous potential for sequence modeling, presenting a compelling avenue for addressing the challenge effectively. Inspired by this, we propose a novel generic and efficient framework (SSD4Rec) for sequential recommendations, which explores the seamless adaptation of Mamba for recommendations. Specifically, SSD4Rec marks the variable- and long-length item sequences with sequence registers and processes the item representations with a novel Masked Bidirectional Structured State Space Duality block. This not only allows for hardware-aware matrix multiplication but also empowers outstanding capabilities in variable-length and long-range sequence modeling. Extensive evaluations on four benchmark datasets demonstrate that the proposed model achieves state-of-the-art performance while maintaining near-linear scalability with user sequence length.
Sequential Recommendation, Recommender Systems, Mamba, State Space Model (SSM), State Space Duality (SSD)
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2024; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
However, directly applying Mamba to sequential recommendation remains highly challenging, primarily due to two key hurdles. Due to varying lengths of users’ historical sequences towards items, most existing sequential modeling methods need padding or truncation operations for sequences with different lengths in recommender systems, resulting in excessive computational overhead (padding) or the potential loss of critical information that could enhance recommendations (truncation), as illustrated in Figure 2. Thus, the first challenge is how to model the dynamic users’ preference with variable- and lone-length historical interaction sequences. Meanwhile, naive Mamba-based models employ a left-to-right uni-directional modeling paradigm, which limits the power of hidden representations for items in historical sequences (Sun et al., 2019). Specifically, each item can only encode information from preceding items, making it difficult to capture precise user preferences and achieve optimal performance in sequential recommendations. This is due to the fact that users’ historical interactions might not follow a rigidly ordered sequence in practice. Thus, the second challenge is how to capture context from both preceding items and subsequent items within the users’ interaction history.
To address such challenges, in this paper, a Structured State-Space Duality-empowered recommendation framework (SSD4Rec) is proposed based on the advanced Mamba architecture to capture the dynamics underlying variable- and long-length user interaction sequences effectively and efficiently. A novel strategy is introduced to encode the user’s sequential interactions into a variable-length item sequence without requiring padding and truncation, thereby accommodating an unlimited input window. Meanwhile, a Masked Bidirectional State Space Duality (Bi-SSD) block based on the Mamba architecture is designed to swallow the numerous item representations and predict corresponding user preferences for proper recommendations.
@@
  • We propose a novel strategy to encode the users’ historical behavior with varying lengths and a Masked Bidirectional Structured State-Space Duality (Bi-SSD) block to capture contextual information, particularly the long-range dependencies, for user behavior sequence modeling.
  • We conducted extensive experiments on four commonly used datasets to empirically showcase the efficacy of SSD4Rec, including its superior recommendation performance and its scalability in predicting the preferences of users with variable- and long-length interaction histories. Specifically, SSD4Rec is 68.47% and 116.67% faster than the representative attention-based and SSM-based methods (i.e., SASRec (Kang and McAuley, 2018) and Mamba4Rec (Liu et al., 2024a)), respectively, when training to extract features on user interactions at the length of 200 with hidden state dimension of 256.
@@ 2.1. Notations and Definitions
The primary goal of sequential recommendation is to predict the next interaction within a sequence. Let $\mathcal{U}=\{u_{1},u_{2},...,u_{|\mathcal{U}|}\}$ and $\mathcal{V}=\{v_{1},v_{2},...,v_{|\mathcal{V}}|\}$ be the sets of users and items, respectively. Each user $u_{k}\in\mathcal{U}$ is characterised by a chronologically ordered sequence $\mathcal{S}_{k}=[v^{u_{k}}_{1},...,v^{u_{k}}_{t},...,v^{u_{k}}_{L_{k}}]$, where $v^{u_{k}}_{t}\in\mathcal{V}$ represents the item that $u_{k}$ has interacted with at time step $t$, and $L_{k}>1$ denotes the interaction sequence length for user $u_{k}$, which may vary from user to user. Notably, for typical sequential recommendation methods, the input sequence length will be uniformed into a fixed hyper-parameter $L$ using the special padding token and truncation. Specifically, if $L_{k}<L$, the sequence is modified to $\mathcal{S}_{k}^{L}=[pad,...,pad,v^{u_{k}}_{1},...,v^{u_{k}}_{L_{k}}]$, with the number of padding instances equal to $(L-L_{k})$. In contrast, if $L_{k}>L$, the sequence is truncated into $\mathcal{S}_{k}^{L}=[v^{u_{k}}_{(L_{k}-L)},...,v^{u_{k}}_{L_{k}}]$. To represent items efficiently, it is common practice to use an embedding table $\mathcal{Q}=\{\mathbf{q}_{1},\mathbf{q}_{2},...,\mathbf{q}_{|\mathcal{V}|}\}$ for the item base $\mathcal{V}$, where $\mathbf{q}_{j}$ is a latent dense vector with dimensionality $D$ for item $v_{j}$.
@@
where $\hat{\mathbf{p}}_{k}\in\mathbb{R}^{D}$ denotes the preference representation of user $u_{k}$, and $r_{k}$ is a score distribution of user $u_{k}$ over all items in the item base $\mathcal{Q}$. This distribution is used to generate a ranked list of items representing the most likely interactions for user $u_{k}$ at step $L+1$.
@@ 2.2. An Overview of the Proposed SSD4Rec
Mamba models have recently demonstrated significant achievements across diverse domains, owing to their exceptional ability to capture intricate dynamics in long sequence modeling (Gu and Dao, 2023; Dao and Gu, 2024). By introducing advanced Structured State Space Models with hardware-aware expansions, Mamba facilitates attention-like modeling capabilities while maintaining a linear complexity in computing. Motivated by these advancements, this paper proposes to explore Mamba’s potential for effective and efficient sequential recommendations. Specifically, unlike the conventional sequential recommendation methods based on RNN or Transformer architecture, the proposed SSD4Rec is designed to handle variable-length item sequences seamlessly without requiring padding or truncation, thereby ensuring that the model receives complete information and retains essential knowledge. As shown in Figure 1, the proposed model consists of two key modules, namely masked variable-length item sequence construction and Bidirectional Structured State Space Duality (Bi-SSD) Layer for user modeling. The first module aims to tackle the core challenge of managing variable-length sequences from diverse users within a certain batch. Meanwhile, the proposed Bi-SSD layer enables attention-like content-aware interaction modeling while upholding linear complexity in computation.
@@
2.3. Masked Variable-length Item Sequence Construction
@@
2.3.1. Masking Operation
It is well-proved that there are frequently unreliable user-item interactions (Fan et al., 2022), especially in long user behaviors, which can impede user modeling in existing data-driven models. To enhance the robustness of SSD4Rec against such noise, we propose to incorporate masking operations into interaction lists during the training phase. This approach creates a challenging task that helps the model develop a comprehensive understanding and generalize effectively. To be specific, we introduce an item-wise masking strategy $\gamma$ following Bernoulli distribution as follows:
(3) $\gamma\sim{\rm Bernoulli}(\rho),$
where $\rho$ is the masking ratio. The Bernoulli distribution is the discrete probability distribution of a random variable, which takes the value 1 with probability $\rho$ and the value 0 with probability $(1-\rho)$. Given the initial representations $\mathcal{E}_{u_{k}}=\{\mathbf{q}_{1},\mathbf{q}_{2},...,\mathbf{q}_{L_{k}}\}^{u_{k}}$ of all the interacted items for user $u_{k}$, we replace the specific items determined by $\gamma$ with a masked embedding $\mathbf{m}$, as shown in Figure 1. This masking operation randomly generates multifarious samples at each training epoch so as to enhance the generalization capability in the proposed model.
2.3.2. Variable-length Item Sequence
@@
(4) $\mathcal{E}_{b}=[\underbrace{\mathcal{E}^{\prime}_{u_{1}},...,\mathcal{E}^{\prime}_{u_{k}},...,\mathcal{E}^{\prime}_{u_{K}}}_{\text{$b$-th batch with size $K$}}],$
where $\mathcal{E}^{\prime}_{u_{k}}\in\mathbb{R}^{L_{k}\times D}$ denotes the masked sequence of item embeddings $\{\mathbf{q}_{1},...,\mathbf{q}_{L_{k}}\}^{u_{k}}$ for user $u_{k}$ with the sequence length of $L_{k}$, and $\mathcal{E}_{b}\in\mathbb{R}^{1\times(\sum L)\times d}$ represents the integrated embedding sequence in the $b$-th batch. For simplicity, we omit the mark ($\prime$); that is, we still use $\mathcal{E}_{u_{k}}$ to denote the masked sequence embeddings for user $u_{k}$. To delineate separate interactions, we incorporate segment registers based on user identifiers (IDs) with the integrated batch input. The registers can be written by
@@ 2.4. Bidirectional State Space Duality (Bi-SSD) for User Modeling
2.4.1. Structured State Space Duality Block
@@
(9) $\mathcal{F}(\mathcal{X})=\textbf{SSM}(\mathbf{A},\mathbf{B},\mathbf{C})(\mathcal{X})$ $=(\mathbf{M}_{1}\mathcal{E}_{u_{1}},...,\mathbf{M}_{k}\mathcal{E}_{u_{k}},...,\mathbf{M}_{K}\mathcal{E}_{u_{K}}),$
@@
where $\mathcal{F}(\mathcal{X})$ is the intermediate output of an SSD block. $\mathbf{M}_{k}$ and $\mathcal{E}_{u_{k}}$ represent the structured matrix and item sequence for user $u_{k}$. By adopting this approach, SSD4Rec implements batch processing of user interaction sequences with varying lengths.
2.4.2. Bi-SSD Layer
@@
By leveraging the state space duality (SSD) block as the foundational component, we construct a network architecture tailored to address the sequential recommendation task, as illustrated in Figure 1. First, in addition to the standard sequential information concerning user behaviors, we propose integrating reverse insights from the interaction history by constructing a bidirectional architecture on top of the unidirectional SSD block. This bidirectional design encourages the model to learn to reason in a reverse manner, potentially capturing more valuable patterns within the interaction histories. Second, the utilization of SSD blocks enables our model to facilitate hardware-aware computation, i.e., block-based matrix multiplication (Dao and Gu, 2024), thereby achieving attention-like content-aware modeling while preserving efficient learning. Moreover, by stacking multiple bidirectional SSD layers, the proposed model can achieve a deeper network structure with an increased number of learnable parameters. Layer-wise normalization and embedding dropout are further utilized between the stacked cells to improve the model’s robustness and stabilize the learning process. Finally, the proposed Bi-SSD layers predict the representations of user preferences in a specific batch, which are used to calculate the ranking score distributions for Top-K sequential recommendations. Given a batch of masked item sequence inputs $\mathcal{X}=(\mathcal{E}_{b},\mathcal{R}_{b})$, the computation procedures of our Bi-SSD layer can be expressed as follows:
  • Forward SSD Block. As detailed in Equation (9), we utilize SSD block to process the masked item sequence input $\mathcal{X}=(\mathcal{E}_{b},\mathcal{R}_{b})$ as follows:
    (11) $\mathcal{F}_{\rm forward}(\mathcal{X})=\textbf{SSM}(\mathbf{A},\mathbf{B},\mathbf{C})(\mathcal{X})$ $=(\mathbf{M}_{1}\mathcal{E}_{u_{1}},...,\mathbf{M}_{k}\mathcal{E}_{u_{k}},...,\mathbf{M}_{K}\mathcal{E}_{u_{K}}).$
  • Backward SSD Block. Before passing the input through the Backward SSD block, we first introduce the user-wise flipping function, which flips the entire masked input sequence from $\mathcal{X}=(\mathcal{E}_{b},\mathcal{R}_{b})$ to $\mathcal{X}^{f}=(\mathcal{E}^{f}_{b},\mathcal{R}_{b})$:
    (12) $\mathcal{E}_{b}=[\mathcal{E}_{u_{1}},...,\mathcal{E}_{u_{k}},...,\mathcal{E}_{u_{K}}]\rightarrow\mathcal{E}^{f}_{b}=[\underbrace{\mathcal{E}^{f}_{u_{1}},...,\mathcal{E}^{f}_{u_{k}},...,\mathcal{E}^{f}_{u_{K}}}_{\text{$b$-th batch with size $K$}}],$
    where $\mathcal{E}^{f}_{u_{k}}$ denotes the flipped user-item sequence corresponding for user $u_{k}$. The flipped input $\mathcal{X}^{f}$ will be processed by the Backward SSD Block as follows:
    (13) $\mathcal{F}_{\rm backward}(\mathcal{X})=\textbf{SSM}(\mathbf{A},\mathbf{B},\mathbf{C})(\mathcal{X}^{f})$ $=(\mathbf{M}_{1}\mathcal{E}^{f}_{u_{1}},...,\mathbf{M}_{k}\mathcal{E}^{f}_{u_{k}},...,\mathbf{M}_{K}\mathcal{E}^{f}_{u_{K}}).$
    Notably, to conserve computational resources and GPU memory usage, we can utilize a shared SSD block for the forward and backward modeling processes, which means the state matrices for the forward SSD block and the backward SSD block are identical.
  • Add & Norm. & Residual Connection Instead of merely summing the output from these three paths, we introduce a weighted indicator $\beta$ to regulate the integration of messages between the forward and backward sequence modeling. As we may stack the Bi-SSD layers, it becomes important to utilize residual connections, which allow the embedding of the last visited item to be propagated to the final layer. Finally, to improve robustness and prevent overfitting, a layer normalization is applied, followed by dropout:
    (14) $\mathcal{F}^{\prime}(\mathcal{X})$ $=\mathcal{F}_{\rm forward}(\mathcal{X})+\beta*\mathcal{F}_{\rm backward}(\mathcal{X})+\mathcal{X},$
    (15) $\mathcal{G}$ $=\text{Dropout}(\text{LayerNorm}(\mathcal{F}^{\prime}(\mathcal{X}))),$
    where $\mathcal{G}=(\hat{\mathcal{E}}_{u_{1}},...,\hat{\mathcal{E}}_{u_{k}},...,\hat{\mathcal{E}}_{u_{K}})$ represents the $K$ hidden item sequences processed by forward and backward SSD blocks, and $\beta$ is a hyper-parameter to balance the contribution of backward sequence modeling.
  • Position-wise Feed-Forward Network. To decode the modeling of user actions in the hidden dimensions, we leverage a position-wise feed-forward network (PFFN) (Liu et al., 2024a) consisting of two dense layers:
    (16) $\text{PFFN}(\mathcal{G})=\text{Dropout}(\text{GELU}(\mathcal{G}\boldsymbol{W}^{(1)}+\boldsymbol{b}^{(1)}))\boldsymbol{W}^{(2)}+\boldsymbol{b}^{(2)},$
    where $W^{(1)}\in\mathbb{R}^{D\times 4D}$, $W^{(2)}\in\mathbb{R}^{4D\times D}$, $b^{(1)}\in\mathbb{R}^{4D}$, $b^{(2)}\in\mathbb{R}^{D}$ are the weights and biases of the two layers, respectively. We incorporate dropout between the two dense layers in the PFFN while the activation function is set to be GELU (Hendrycks and Gimpel, 2016). PFFN can be considered a channel-mixing process that enhances the model’s ability to understand and process complex patterns by mixing the information from each channel that passes through the output of bidirectional SSD blocks. Furthermore, to enhance robustness and prevent overfitting, we incorporate residual connection, alongside dropout and layer normalization, to the PFFN. The final output of each Bi-SSD layer $\mathcal{H}$ can be calculated by:
    (17) $\mathcal{H}=\text{Dropout}(\text{LayerNorm}(\text{PFFN}(\mathcal{G})+\mathcal{G})).$
  • Stacking Bi-SSD Layers. As detailed above, user preferences throughout the entire user-item sequence can be easily captured by a single Bi-SSD layer. Furthermore, Bi-SSD layers can be stacked to allow the model to capture more complex user preferences. However, adding more layers to the Bi-SSD increases training time due to the higher number of model parameters. We will analyze the appropriate number of Bi-SSD layers to stack in the later experimental section. In summary, SSD4Rec refines the hidden representations of $m$-th layer as follows:
[…]
    (20) $\mathcal{G}^{m}$ $=\text{Dropout}(\text{LayerNorm}(\mathcal{F}^{\prime}(\mathcal{H}^{m}))),$
    (21) $\mathcal{F}^{\prime}(\mathcal{H}^{m})$ $=\mathcal{F}_{\rm forward}(\mathcal{H}^{m})+\beta*\mathcal{F}_{\rm backward}(\mathcal{H}^{m})+\mathcal{H}^{m},$
    where $m=0,1,...,M$, and $\mathcal{H}^{0}=\mathcal{X}$.